OpenStudy (anonymous):

pls help what to do when the perimeter of the rhombus is 180 and the length of one of its diagonal is 72cm what is measure of the other diagonal

OpenStudy (gorv):

area = d1*d2*1/2 d1 & d2 are diagonals of rhombus 180=72*d2*1/2=36*d2 d2=180/36=5 cm

OpenStudy (gorv):

@HARSH3166

OpenStudy (anonymous):